# Study of hybrid stars with nonstrange quark matter cores

### v2

@@ II EOS of nonstrange quark matter
Figure 1: The chiral susceptibilities $\chi_{\rm s}$ versus temperature for the three parameter sets defined in Table 1 at $\mu=0$. The three vertical lines denote where the peaks of $\chi_{\rm s}$ are located, i.e., $T_{\rm pc}=165,172,181$ MeV.
@@
Figure 2: The scaled quark condensates versus $T$ for the three parameter sets in Table 1 at $\mu=0$. The corresponding pseudo-critical points are also marked on these curves, respectively.
Fig. 1 plots the chiral susceptibilities versus temperature. We can see that if $G_{1}$ varies from 1.935 to 1.972 GeV-2, $T_{\rm pc}$ will vary from 165 to 181 MeV, thus satisfy the constraint from lattice simulations. Considering Eq. (10) at $T=\mu=0$, the corresponding value of $G_{2}$ can be derived from $-1.582$ to $-0.743$ GeV-5. The whole parameter sets of the modified 2-flavor NJL model in this work are shown in Table 1, where in addition to the two boundary values of $G_{1}$ that meet the requirement of lattice simulations, we also take the parameter set of $G_{1}=1.950$ GeV-5 into account, in order to study the influence of different parameter sets on the EOSs.
@@
$\quad m$ $\,\,\Lambda_{\rm UV}$ $\,\,M_{\rm{u,d}}$ $-\langle\bar{\psi}\psi\rangle^{\frac{1}{3}}$ $\quad\,\,G$ $\quad\,\,\,G_{1}$ $\quad\,\,\,G_{2}$
@@
1.9351.935$\quad$ 1.9351.935 $\,\,\,-1.582$
    3.5  1324    180180\,\,\,180180      353353\,\,\,\,\,353353    2.0052.005\,\,\,2.0052.005 1.9501.950$\quad$ 1.9501.950 $\,\,\,-1.242$
1.9721.972$\quad$ 1.9721.972 $\,\,\,-0.743$
@@
Fig. 2 plots the scaled order parameter of chiral phase transition ($\langle\bar{\psi}\psi\rangle/\langle\bar{\psi}\psi\rangle_{0}$) versus temperature. We can find that $\langle\bar{\psi}\psi\rangle/\langle\bar{\psi}\psi\rangle_{0}$ decreases smoothly from one to zero as temperature increases, thus the transition at $\mu=0$ is the crossover, consistent with the simulation result of lattice QCD. For the three parameter sets in Table. 1, the pseudo-critical temperature increases with the increase of $G_{1}$.
Now we extend our calculation to finite chemical potentials at $T=0$ to get EOSs of the quark matter. After solving Eq. (4) with the modification of Eq. (10), we can get the dependence of effective quark mass $M_{\rm u,d}$ on the chemical potential, which is shown in Fig. 3. It can be seen that the chiral phase transition at $T=0$ is still a crossover for each parameter set in this work. The pseudo-critical chemical potentials $\mu_{\rm pc}$, determined by the chiral susceptibility, are 272, 280, and 293 MeV for the three parameter sets of G1935, G1950 and G1972 defined in Table 1, respectively.
@@
Figure 3: The effective mass of $u,d$ quarks versus chemical potentials $\mu$ at $T=0$ for the parameter sets defined in Table 1. The marked points on the three lines are the pseudo-critical points with $\mu_{\rm pc}=272,280,293$ MeV, respectively.
@@
In the framework of the NJL model, it is demonstrated that whether the first-order chiral phase transition occurs at $T=0$ (when $m\neq 0$) depends on the regularization scheme that is employed Buballa (2005); Zhang et al. (2016); Kohyama et al. (2015). In Ref. Buballa (2005), the three-momentum cutoff regularization is used and a first-order phase transition happens at $T=0$. However, in Ref. Zhang et al. (2016), the authors use the PTR and find a crossover in the phase transition region at $T=0$. Actually, in Ref. Kohyama et al. (2015), it is clarified that the low current quark mass ($m\leq 4$ MeV) can result in a crossover at $T=0$ for both PTR and three-momentum cutoff regularization. It is noted that $m=3.5$ MeV in this work just corresponds to the case of low current quark mass, thus a crossover occurs in the chiral phase transition region.
@@
Figure 4: The number density of $u,d$ quarks versus chemical potentials $\mu$ at $T=0$ for the parameter sets defined in Table 1.
The dependence of quark number density on the chemical potential at $T=0$ is shown in Fig. 4. We can see that for the three parameter sets in Table 1, the number densities of $u,d$ quarks are only slightly different in the crossover region, and a smaller value of $G_{1}$ corresponds to a larger number density for the same chemical potential.
@@
where $\rho_{\rm e}=\mu_{\rm e}^{3}/3\pi^{2}$ is the number density of electrons at $T=0$. For the three parameter sets defined in Table 1, the relation between the baryon density $\rho_{\rm B}$ and baryon chemical potential $\mu_{\rm B}$ is shown in Fig. 5, where $\mu_{\rm B}=\mu_{\rm u}+2\mu_{\rm d}$ and $\rho_{\rm B}=(\rho_{\rm u}+\rho_{\rm d})/3$. In Fig. 5, the $\rho_{\rm B}-\mu_{\rm B}$ relations in the cases of $G_{1}=1.935,1.950,1.972$ GeV-2 are almost the same except for the crossover region, where a smaller value of $G_{1}$ corresponds to a larger $\rho_{\rm B}$ for the same $\mu_{\rm B}$.
Figure 5: The baryon density versus baryon chemical potential of the quark system for the parameter sets defined in Table 1.
According to the definition, the EOS of dense quark matter at $T=0$ is Zong and Sun (2008)
@@
and the energy density of the quark system can be expressed as Yan et al. (2012); Benvenuto and Lugones (1995)
@@
It is noted that $P(\mu=0)$ in Eq. (14) is irrelevant to the chemical potential. It corresponds to the negative vacuum pressure and cannot be determined in a model-independent way. Just like the vacuum bag constant ($-B$) in the MIT bag model, we treat it as a phenomenological parameter in this work. In general, $B^{1/4}$ is in a range of 100-200 MeV and should be constrained from ground experiments and astronomical observations Lu et al. (1998); Song et al. (1992). In this work, we will constrain its value in light of recent neutron star observations.
III Structure of hybrid stars
@@
where $\mu_{\rm B,c}$ is the critical baryon chemical potential of the hadron-quark phase transition, which is around 1.4 GeV in Fig. 6 for $G_{1}=1.935,1.950,1.972$ GeV-2 and $B^{1/4}=150,153,161$ MeV. Note that the dense hadronic matter in this work is described by the BSK21 parametrization of the Skyrme interaction Goriely et al. (2010); Chamel et al. (2011), which is strongly favored by the nuclear-mass measurements Chamel et al. (2011). The hybrid EOS can be written as
$P(\mu_{\rm B})=\left\{\begin{array}[]{lcl}P_{\rm H},\,\,\,when\,\,\mu_{\rm B}\leq\mu_{\rm B,c},\\ P_{\rm Q},\,\,\,when\,\,\mu_{\rm B}>\mu_{\rm B,c}.\end{array}\right.$ (15)
@@
$\epsilon(\mu_{\rm B})=\left\{\begin{array}[]{lcl}\epsilon_{\rm H},\,\,\,when\,\,\mu_{\rm B}\leq\mu_{\rm B,c},\\ \epsilon_{\rm Q},\,\,\,when\,\,\mu_{\rm B}>\mu_{\rm B,c},\end{array}\right.$ (16)
@@
Because $G_{1}$ has already been constrained in a range of (1.935, 1.972) GeV-2 by lattice simulations in Sec. II, the other adjustable parameter $B^{1/4}$ in hybrid EOSs can be constrained by neutron star observations. In fact, considering the 2 $M_{\odot}$ constraint, the lower limits of $B^{1/4}$ in the cases of $G_{1}=1.935,1.950,1.972$ GeV-2 should be 150, 149, and 148 MeV, respectively, while the stability of hybrid stars with a quark matter core yields the upper limits of $B^{1/4}$ as 161, 158, and 157 MeV, respectively. In addition, we will also choose $B^{1/4}=153$ MeV for comparison in the following calculation to study the influence of $B^{1/4}$ on the EOSs and on the structure of hybrid stars.
In Fig. 6, the influences of $G_{1}$ and $B^{1/4}$ on the pressure of the quark matter are presented in the subgraph (a) and (b) with $B^{1/4}=153$ MeV and $G_{1}=1.935$ GeV-2, respectively. When $B^{1/4}=153$ MeV, there is little difference between the pressures of the quark matter as long as $G_{1}\in(1.935,1.972)$ GeV-2. However, when $G_{1}=1.935$ GeV-2, the pressure difference of the quark matter with $B^{1/4}=150,153,161$ MeV is obvious, and a smaller $B^{1/4}$ leads to a higher $P$ for the same $\mu_{\rm B}$.
@@
Figure 6: The pressure versus baryon chemical potential ($\mu_{\rm B}$) for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2. The pressure of the hadronic matter described by BSK21 is also presented for a comparison. G1935, G1950, G1972 refer to $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and B150, B153, B161 refer to $B^{1/4}=150,153,161$ MeV, respectively.
@@
In Fig. 7, we present the $\epsilon-P$ relations of the hadronic matter, the quark matter and hybrid EOSs with the Maxwell construction. Each point marked with “x” represents the critical point of the corresponding first-order phase transition, which is denoted with ($P_{pt}$, $\epsilon_{pt}$) in the following, and the other marked point on each hybrid EOS refers to the center of the most massive hybrid star, which is denoted with ($P_{\rm c}$, $\epsilon_{\rm c}$) in the following. When $B^{1/4}=153$ MeV, the locations of ($P_{pt}$, $\epsilon_{pt}$) for $G_{1}=1.935,1.950$ GeV-2 are the same. However, due to the small but non-negligible differences of the hybrid EOSs in these two cases, the corresponding ($P_{\rm c}$, $\epsilon_{\rm c}$) are different. When $G_{1}=1.935$ GeV-2, the differences of both ($P_{\rm pt}$, $\epsilon_{\rm pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) are obvious in the cases of $B^{1/4}=150,153,161$ MeV, and a larger $B^{1/4}$ will lead to larger values of ($P_{pt}$, $\epsilon_{pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$). The corresponding ($P_{pt}$, $\epsilon_{pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) points of nine representative hybrid EOSs are listed in Table 2.
@@
Table 2: The corresponding ($P_{pt}$, $\epsilon_{pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) points of nine representative hybrid EOSs.
$\quad\,G_{1}$ $\quad B^{1/4}$ $\,\,\,\,\,(P_{pt},\epsilon_{pt})$ $\,\,\,\,\,\,\,(P_{\rm c},\epsilon_{\rm c})$
$[{\rm GeV}^{-2}]$ $\,\,\,[{\rm MeV}]$ $\,[{\rm MeV}\cdot{\rm fm}^{-3}]$ $\,[{\rm MeV}\cdot{\rm fm}^{-3}]$
150150$\quad$ 150150 $(164.0,634.1)$ $(194.6,957.7)$
    1.935 153153$\quad$ 153153 $(180.7,662.1)$ $(201.0,998.9)$
161161$\quad$ 161161 $(219.9,725.0)$ $(230.5,1153.0)$
149149$\quad$ 149149 $(164.0,634.1)$ $(193.6,953.6)$
    1.950 153153$\quad$ 153153 $(180.7,662.1)$ $(193.5,982.2)$
158158$\quad$ 158158 $(209.3,708.4)$ $(220.9,1104.6)$
148148$\quad$ 148148 $(164.0,634.1)$ $(194.8,958.9)$
    1.972 153153$\quad$ 153153 $(189.7,677.0)$ $(200.2,1011.3)$
157157$\quad$ 157157 $(209.3,708.4)$ $(220.1,1102.8)$
@@
Figure 7: The $\epsilon-P$ relation of the quark matter and hybrid EOS when (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2. The $\epsilon-P$ relation of the hadronic matter described by BSK21 is shown by the red line. G1935, G1950, G1972 (hEOS1935, hEOS1950, hEOS1972) refers to the quark matter (hybrid EOS) when $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and B150, B153, B161 (hEOS150, hEOS153, hEOS161) refers to quark matter (hybrid EOS) with $B^{1/4}=150,153,161$ MeV, respectively.
@@
Once the EOS is determined, we can solve the Tolman-Oppenheimer-Volkoff (TOV) equation numerically to get the $M-R$ and mass-central energy density ($M-\epsilon_{\rm c}$) relations. In Fig. 8 and Fig. 9, to study the influence of $G_{1}$ and $B^{1/4}$ on the $M-R$ and $M-\epsilon_{\rm c}$ relations, the corresponding results for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2 are shown, respectively. In Fig. 8, the most massive quark star with $G_{1}=1.935,1.950,1.972$ GeV-2 ($B^{1/4}=150,153,161$ MeV) is about 1.52 $M_{\odot}$ (1.57 $M_{\odot}$), not satisfying the 2 $M_{\odot}$ constraint. In addition, the quark stars described by the modified NJL model in this work cannot fulfill the recent $M-R$ constraint from the NICER mission (PSR J0740+6620 Miller et al. (2021) and PSR J0030+0451 Miller et al. (2019a)). However, the hybrid EOSs obtained with the Maxwell construction approach in this work can produce hybrid stars in consistent with these astronomical observations, although their quark matter cores are relatively small (about 0.02 $M_{\odot}$).
Figure 8: The $M-R$ relations of quark stars and hybrid stars for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2. The $M-R$ relation of the hadron star described by BSK21 is shown by the red line. QS1935, QS1950, QS1972 (HS1935, HS1950, HS1972) refer to quark stars (hybrid stars) with $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and QS150, QS153, QS161 (HS150, HS153, HS161) refer to quark stars (hybrid stars) with $B^{1/4}=150,153,161$ MeV, respectively. The constraints from the NICER mission (PSR J0740+6620 Miller et al. (2021) and PSR J0030+0451 Miller et al. (2019a)) are shown as dark (light) colored areas, referring to the $68\%$ ($95\%$) confidence levels.

### v3

@@ 2 EOS of nonstrange quark matter
Figure 1: The chiral susceptibilities $\chi_{\rm s}$ versus temperature for different $G_{1}$. $m$ is fixed as 7.5 MeV. The three vertical lines denote the peak place of $\chi_{\rm s}$, i.e., $T_{\rm pc}=165,181,247.1$ MeV.
@@
Figure 2: The scaled quark condensates versus $T$ for different $G_{1}$. $m$ is fixed as 7.5 MeV. The corresponding pseudo-critical points are also marked on these curves, respectively.
According to our calculations, when the current quark mass $m$ is in the range of 3.5 – 10 MeV, $T_{\rm pc}$ will be larger than 181 MeV in the conventional NJL model. Generally, $T_{\rm pc}$ will decrease as $G_{1}$ decreases. In Fig. 1, the chiral susceptibility for different $G_{1}$ is plotted, taking $\mu=0$ and $m=7.5$ MeV. As $G_{1}$ varies from 4.358 to 4.552 GeV-2, $T_{\rm pc}$ changes from 165 to 181 MeV, satisfying the constraint from lattice simulations. However, when $G_{1}=G=5.29$ GeV-2, $T_{\rm pc}$ in the conventional NJL model is 247.1 MeV, which is much larger than the corresponding value of lattice simulations.
The QCD sum rule at the renormalization scale of 1 GeV suggests the $u$ quark condensate should be $229\pm 33$ MeV3  (Dosch & Narison, 1998) or $242\pm 15$ MeV3  (Jamin, 2002). Considering this constraint, the whole parameter sets of the modified 2-flavor NJL model in this work are shown in Table 1. For a certain current quark mass, we take two boundary values for $G_{1}$, corresponding to the cases with $T_{\rm pc}=165$ and 181 MeV, respectively. Similar to previous studies under the proper-time regularization (Klevansky, 1992; Kohyama et al., 2015; Li et al., 2017), the effective quark masses in Table 1 are in a range of 190 – 250 MeV, which are smaller than 300 MeV. As a comparison, it has been shown that for the same current quark mass, the three-momentum cutoff regularization leads to a larger effective quark mass and a smaller momentum cutoff  (Klevansky, 1992; Ratti et al., 2007; Kohyama et al., 2015, 2016).
@@
$\quad m$ $\,\,\Lambda_{\rm UV}$ $\,\,M$ ${-}\langle\bar{u}u\rangle^{\frac{1}{3}}$ $\quad\,\,G$ $\quad\,\,\,G_{1}$ $\quad\,\,\,G_{2}$
@@
    4.5  1151    191191\,\,\,191191      258258\,\,\,\,\,258258    2.7272.727\,\,\,2.7272.727 2.5632.563$\quad$ 2.5632.563 $\,\,\,-4.791$
    4.5  1151    191191\,\,\,191191      258258\,\,\,\,\,258258    2.7272.727\,\,\,2.7272.727 2.6272.627$\quad$ 2.6272.627 $\,\,\,-2.924$
    5.8  1002    206206\,\,\,206206      237237\,\,\,\,\,237237    3.7613.761\,\,\,3.7613.761 3.3753.375$\quad$ 3.3753.375 $\,\,\,-14.50$
    5.8  1002    206206\,\,\,206206      237237\,\,\,\,\,237237    3.7613.761\,\,\,3.7613.761 3.4853.485$\quad$ 3.4853.485 $\,\,\,-10.36$
    7.5  873    225225\,\,\,225225      217217\,\,\,\,\,217217    5.2905.290\,\,\,5.2905.290 4.3584.358$\quad$ 4.3584.358 $\,\,\,-45.29$
    7.5  873    225225\,\,\,225225      217217\,\,\,\,\,217217    5.2905.290\,\,\,5.2905.290 4.5524.552$\quad$ 4.5524.552 $\,\,\,-35.86$
    9.5  774    250250\,\,\,250250      201201\,\,\,\,\,201201    7.3997.399\,\,\,7.3997.399 5.2265.226$\quad$ 5.2265.226 $\,\,\,-133.8$
    9.5  774    250250\,\,\,250250      201201\,\,\,\,\,201201    7.3997.399\,\,\,7.3997.399 5.5725.572$\quad$ 5.5725.572 $\,\,\,-112.5$
@@
Fig. 2 plots the scaled order parameter of chiral phase transition ($\langle\bar{\psi}\psi\rangle/\langle\bar{\psi}\psi\rangle_{0}$) versus temperature when $m=7.5$ MeV. We can find that $\langle\bar{\psi}\psi\rangle/\langle\bar{\psi}\psi\rangle_{0}$ decreases smoothly from one to zero as temperature increases, thus the transition at $\mu=0$ is the crossover, consistent with the simulation result of lattice QCD.
Now we extend our calculation to finite chemical potentials at $T=0$ to get EOSs of the quark matter. After solving Eq. (6) with the modification of Eq. (3), we can get the dependence of effective quark mass $M$ on the chemical potential, which is shown in Fig. 3. It can be seen that for the dense quark systems with a fixed $T_{\rm pc}$, a larger $m$ leads to a larger effective quark mass in the vacuum, and thus a larger gap will emerge when the chiral phase transition occurs. Specifically, when $T_{\rm pc}=181$ MeV, the crossover occurs for the systems with $m<9$ MeV, and the first order phase transition occurs for the systems with $m>9$ MeV. The critical chemical potential ($\mu_{\rm c}$) is around 285 MeV. As a comparison, we also present the results for a fixed $m$ but with different values of $G_{1}$ in Fig. 3. As $G_{1}$ increases, $\mu_{\rm c}$ will also increase. Note that the green dashed line in Fig. 3 corresponds to normal NJL model, in which the pseudo-critical chemical potential is about 350 MeV. It is much larger than the corresponding results of those cases with $G_{1}\in(4.430,4.552)$ GeV-2.
@@
Figure 3: The effective mass of quarks versus chemical potentials $\mu$ at $T=0$. Note that m58, m75, m95 correspond to three cases with $m=5.8$, 7.5, 9.5 MeV, respectively, in which we have taken $T_{\rm pc}=181$ MeV. Similarly, in the three cases of m75G5290, m75G4552 and m75G4430, we have taken $m=7.5$ MeV and $G_{1}=5.290$, 4.552, 4.43 GeV-2, respectively. The (pseudo-)critical points are marked on the lines correspondingly.
@@
In the framework of the NJL model, it is demonstrated that whether the first-order chiral phase transition occurs at $T=0$ (when $m\neq 0$) depends on the regularization scheme that is employed  (Buballa, 2005; Zhang et al., 2016; Kohyama et al., 2015). In Buballa (2005), the three-momentum cutoff regularization is used and a first-order phase transition happens at $T=0$. However, in Zhang et al. (2016), the authors use the PTR and find a crossover in the phase transition region at $T=0$. Actually, in Kohyama et al. (2015), it is clarified that the low current quark mass ($m\leq 4$ MeV) can result in a crossover at $T=0$ for both PTR and three-momentum cutoff regularization.
According to lattice QCD simulations, we notice that at zero temperature and under the PTR, when $m>7.5$ MeV, $\mu_{\rm c}$ of the modified NJL model is significantly lower than that in the normal NJL model. Sometimes even a first-order phase transition occurs. For example, from the phase diagram of the normal NJL model derived by Kohyama et al. (2015), we see that the crossover can only occur at $T=0$ when $m<10$ MeV. However, according to our calculations, a first-order phase transition has already occurred when $m>9$ MeV in the modified NJL model.
@@
Figure 4: The number density of $u,d$ quarks versus chemical potentials $\mu$ at $T=0$. Line styles are the same as those in Fig. 3.
The dependence of quark number density on the chemical potential at $T=0$ is shown in Fig. 4. We can see that a larger $m$ leads to a lower quark number density when $T_{\rm pc}=181$ MeV and $\mu\in(220-290)$ MeV. For a fixed $m$ with different values of $G_{1}$, the quark number densities are only slightly different in the crossover region according to the lattice QCD simulation results on $T_{\rm pc}$. However, they are quite different in normal NJL models.
@@
where $\rho_{\rm e}=\mu_{\rm e}^{3}/3\pi^{2}$ is the number density of electrons at $T=0$. The relation between the baryon density $\rho_{\rm B}$ and baryon chemical potential $\mu_{\rm B}$ is shown in Fig. 5, where $\mu_{\rm B}=\mu_{\rm u}+2\mu_{\rm d}$ and $\rho_{\rm B}=(\rho_{\rm u}+\rho_{\rm d})/3$. In Fig. 5, the shape of each curve is quite similar to that in Fig. 4. Interestingly, in the modified NJL model, the $\rho_{\rm B}-\mu_{\rm B}$ relation of the quark system is almost identical even when different parameter sets are adopted. The difference mainly occurs in the phase transition region.
Figure 5: The baryon density versus baryon chemical potential of the quark system. Line styles are the same as those in Fig. 3.
According to the definition, the EOS of dense quark matter at $T=0$ is (Zong & Sun, 2008)
@@
and the energy density of the quark system can be expressed as (Yan et al., 2012; Benvenuto & Lugones, 1995)
@@
Note that $P(\mu=0)$ in Eq. (12) is irrelevant to the chemical potential. It represents the negative vacuum pressure, which corresponds to the vacuum bag constant ($-B$) in the MIT bag model. In the NJL model, it can be calculated by using Eq. (2.60) of Buballa (2005) under the mean field approximation. In this way, the bag constant is regarded as the pressure difference between the trivial and the non-trivial vacuum. It is known that the trivial and non-trivial vacuum can be described by the (quasi-)Wigner and Nambu solution of the gap equation, respectively. As suggested by Xu et al. (2018); Cui et al. (2018); Li et al. (2019a); Wang et al. (2019), the vacuum pressure is defined as
$P(\mu=0)$ $=$ $P(M_{N})-P(M_{W})$ (14)
$=$ $\Omega(0,0,\langle\bar{\psi}\psi\rangle_{W})-\Omega(0,0,\langle\bar{\psi}\psi\rangle_{N}),$
where $M_{N}$ and $M_{W}$ represents the Nambu and quasi-Wigner solution of the gap equation, respectively, and $\langle\bar{\psi}\psi\rangle_{N}$ and $\langle\bar{\psi}\psi\rangle_{W}$ are the corresponding quark condensates.
3 Structure of hybrid stars
@@
where $\mu_{\rm B,c}$ is the critical baryon chemical potential of the hadron-quark phase transition, which is around 1.5 and 1.6 GeV in the modified NJL model and the normal NJL model, respectively. Note that the dense hadronic matter in this work is described by the APR EOS with $A18+\delta\nu+UIX^{\ast}$ interaction (Akmal et al., 1998), which is strongly favored by recent neutron star observations. The hybrid EOS can be written as
$P(\mu_{\rm B})=\left\{\begin{array}[]{lcl}P_{\rm H},\,\,\,when\,\,\mu_{\rm B}\leq\mu_{\rm B,c},\\ P_{\rm Q},\,\,\,when\,\,\mu_{\rm B}\geq\mu_{\rm B,c}.\end{array}\right.$ (16)
@@
$\epsilon(\mu_{\rm B})=\left\{\begin{array}[]{lcl}\epsilon_{\rm H},\,\,\,when\,\,\mu_{\rm B}\leq\mu_{\rm B,c},\\ \epsilon_{\rm Q},\,\,\,when\,\,\mu_{\rm B}\geq\mu_{\rm B,c},\end{array}\right.$ (17)
@@
In Sec. 2, the parameter sets have been preliminarily constrained by lattice simulations, and the results are shown in Table 1. Here they will be further constrained based on recent measurements on neutron star mass, radius, and tidal deformability. According to our calculations, a 2 $M_{\odot}$ neutron star can provide a strong constraint on the lower limit of $G_{1}$. It rules out the parameter sets with $m<5.8$ MeV and $m>9.5$ MeV. On one hand, when $m<5.8$ MeV, the most massive hybrid star mass cannot reach 2 $M_{\odot}$. On the other hand, when $m>9.5$ MeV, the pressure of quark matter is higher than that of hadronic matter, and the most massive quark star is also lighter than 2 $M_{\odot}$.
In the following, to study the influence of various parameters on the EOS and on the structure of hybrid stars, we choose five representative quark EOSs and their corresponding hybrid EOSs for comparison. The corresponding parameter sets are $m=5.8$ MeV with $G_{1}=3.485$ GeV-2, $m=7.5$ MeV with $G_{1}=4.430,4.552,5.290$ GeV-2, respectively, and $m=9.5$ MeV with $G_{1}=5.572$ GeV-2. Specifically, the parameter sets with $(m,G_{1})=$(5.8 MeV, 3.485 GeV-2), (7.5 MeV, 4.552 GeV-2), (9.5 MeV, 5.572 GeV-2) correspond to the case of $T_{\rm pc}=181$ MeV, from which we can see the influence of $m$ on the results. The parameter sets with $(m,G_{1})=$(7.5 MeV, 4.430 GeV-2), (7.5 MeV, 4.552 GeV-2), (7.5 MeV, 5.290 GeV-2) correspond to the case of $m=7.5$ MeV, from which we can find the influence of $G_{1}$ on the results.
According to the discussion on the vacuum pressure $P(\mu=0)$ in Eq. 14, the bag constant $B^{1/4}$ corresponding to the parameter sets of $(m,G_{1})$=(5.8 MeV, 3.485 GeV-2), (7.5 MeV, 4.430 GeV-2), (7.5 MeV, 4.552 GeV-2), (9.5 MeV, 5.572 GeV-2) can be calculated as 122.7 MeV, 120.9 MeV, 123.6 MeV, 125.4 MeV, respectively. Comparing with the normal NJL model with $(m,G_{1})$ = (7.5 MeV, 5.290 GeV-2) and $B^{1/4}=137.1$ MeV, the modified NJL model under the constraint of $T_{\rm pc}$ has a smaller $B^{1/4}$, which ranges in 120 – 125 MeV. For a fixed $T_{\rm pc}$, a larger $m$ will lead to a slightly larger $B^{1/4}$. Note that the range of $B$ in this work is consistent with that in  Song et al. (1992). Considering the medium effect, the study  (Lu et al., 1998) also indicated the decline of $B$ in the framework of the quark-meson coupling model.
In Fig. 6, the influences of $m$ and $G_{1}$ on the pressure of the quark matter are presented in the left panel and right panel with $T_{\rm pc}=181$ MeV and $m=7.5$ MeV, respectively. We can see that after considering the constraint of lattice simulation results on $T_{\rm pc}$, the EOSs of quark matter under different parameter sets are very close to each other. It could also be seen that the deconfinement phase transition occurs at about $\mu_{B}\sim 1.5$ GeV. The left panel shows the EOSs corresponding to $T_{\rm pc}=181$ MeV and $m=5.8$, 7.5, 9.5 MeV, respectively. Similarly, the right panel shows the EOSs corresponding to $m=7.5$ MeV and $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively. The EOSs with $G_{1}=4.430$ and 4.552 GeV-2 are very close to each other and their corresponding $T_{\rm pc}$ values are in accordance with the lattice simulation results. However, both of these two EOSs are quite different from that of $G_{1}=5.290$ GeV-2, which corresponds to the case of normal NJL model.
@@
Figure 6: The pressure versus baryon chemical potential ($\mu_{\rm B}$) for $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. The pressure of the hadronic matter described by BSK21 is also presented for a comparison. Here, m58, m75, m95 refer to $m=5.8$, 7.5, 9.5 MeV, respectively, and G5290, G4552, G4430 refer to $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively.
@@
In Fig. 7, we present the $\epsilon-P$ relations of the hadronic matter, the quark matter and hybrid EOSs with the Maxwell construction. Each point marked with “x” represents the critical point of the corresponding first-order phase transition, which is denoted as ($P_{\rm pt}$, $\epsilon_{\rm pt}$) hereafter. The other marked point on each hybrid EOS refers to the center of the most massive hybrid star. Here we denote it as ($P_{\rm c}$, $\epsilon_{\rm c}$). We can see that according to the constraint of lattice simulation results on $T_{\rm pc}$, $\epsilon_{\rm c}$ ($\epsilon_{\rm pt}$) of the hybrid EOSs only varies slightly among different parameter sets. Generally, a larger $G_{1}$ will lead to a larger value $P_{\rm c}$ and $\epsilon_{\rm c}$. The corresponding ($P_{\rm pt}$, $\epsilon_{\rm pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) points of five representative hybrid EOSs are listed in Table 2.
@@
Table 2: The corresponding ($P_{\rm pt}$, $\epsilon_{\rm pt}$) and ($P_{\rm c}$, $\epsilon_{\rm c}$) points of five representative hybrid EOSs.
$\quad\,m$ $\quad\,\,G_{1}$ $\,\,\,\,\,(P_{pt},\epsilon_{pt})$ $\,\,\,\,\,\,\,(P_{\rm c},\epsilon_{\rm c})$
$\,\,[{\rm MeV}]$ $[{\rm GeV}^{-2}]$ $\,[{\rm MeV}\cdot{\rm fm}^{-3}]$ $\,[{\rm MeV}\cdot{\rm fm}^{-3}]$
   5.8     3.4853.485\,\,\,\,3.4853.485 $(299.9,856.5)$ $(348.6,1312.3)$
    4.4304.430\,\,\,\,4.4304.430 $(299.9,856.5)$ $(306.2,1179.6)$
   7.5     4.5524.552\,\,\,\,4.5524.552 $(299.9,856.5)$ $(332.5,1283.2)$
    5.2905.290\,\,\,\,5.2905.290 $(415.8,980.1)$ $(424.1,1707.7)$
   9.5     5.5725.572\,\,\,\,5.5725.572 $(326.6,886.3)$ $(350.5,1358.2)$
@@
Figure 7: The $\epsilon-P$ relation of the quark matter and hybrid EOS when $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. The $\epsilon-P$ relation of the hadronic matter described by APR is shown by the red line. Here, m58, m75, m95 (APR-m58, APR-m75, APR-m95) refer to the quark matter (hybrid EOS) with $m=5.8$, 7.5, 9.5 MeV, respectively, and G5290, G4552, G4430 (APR-G5290, APR-G4552, APR-G4430) refer to quark matter (hybrid EOS) with $G_{1}=5.290$, 4.552, 4.430 GeV-2, respectively.
@@
Once the EOS is determined, we can solve the Tolman-Oppenheimer-Volkoff (TOV) equation numerically to get the $M-R$ and mass-central energy density ($M-\epsilon_{\rm c}$) relations. In Fig. 8 and Fig. 9, to study the influence of $m$ and $G_{1}$ on the $M-R$ and $M-\epsilon_{\rm c}$ relations, the corresponding results for $T_{\rm pc}=181$ MeV and $m=7.5$ MeV are shown in the left panel and right panel, respectively. In Fig. 8, the most massive quark star for $m=5.8$, 7.5, 9.5 MeV ($G_{1}=5.290$, 4.552, 4.430 GeV-2) is about 1.892 $M_{\odot}$ (1.879 $M_{\odot}$), not satisfying the 2 $M_{\odot}$ constraint. In addition, the quark stars cannot fulfill the recent $M-R$ constraint from the NICER measurement of PSR J0030+0451 (Miller et al., 2019b), although the other constraint from PSR J0740+6620 (Miller et al., 2021) is marginally fulfilled by the quark EOSs with the parameter sets of $(m,G_{1})=(5.8,3.485)$ and $(7.5,4.430)$. However, the hybrid EOSs obtained with the Maxwell construction approach in this work can produce hybrid stars in consistent with these astronomical observations, although their quark matter cores are relatively small (about 0.03 $M_{\odot}$).
Figure 8: The $M-R$ relations of quark stars and hybrid stars for $T_{\rm pc}=181$ MeV (left panel) and $m=7.5$ MeV (right panel), respectively. Line styles are the same as in Fig. 7.